a company produced ethyl el alcohol and claimed to jave a mean alcohol content of 70%.A random sample of 80 of ethyl alcohol bottles was taken as sample to verify this claim.it was found out that the mean alcohol content is 65% with a standard deviation of 2%. test the claim at 5% level of significance and assume that the population is normallt distributed


1
Expert's answer
2021-05-31T16:29:47-0400

Given claim

The mean alcohol content in ethyl alcohol produced by company is of 70%.

The null and alternate hypothesis are shown below

"H_0: μ=70 \\

H_a: μ≠70"

The sample mean alcohol content is "\bar{X}=65\%" and standard deviation is σ=2%

The level of significance is 5%.

The population is normally distributed and as the sample size is high the sample is also treated as normally distributed.

The test statistic is calculated as shown below

"Z= \frac{\bar{X} -\mu}{\sigma} \\

Z= \frac{65-70}{2}=-2.5"

This is a two tailed hypothesis test.

"P-value =2 \times P(Z<-2.5)"

With respect to Z table we get P(Z<-2.5)=0.0062

So p value is calculated as shown below

"p-value =2 \times P(Z< -2.5)=2 \times 0.0062=0.0124"

The p-value 0.0124 is less than the level of significance 0.05 or 5%. Thus we have sufficient evidence to reject the null hypothesis. So we conclude that the alcohol content in ethyl alcohol produced by company is not 70%.